How large is the affected region of the chromosome if λ = 10 − 8 and s = 0 . 1 ?
Answer #1

The chromosomal location of human constant region light chain Ig genes has been determined by analyzing a group of human fibroblast somatic cell hybrids with nucleic acid probes prepared from cloned human kappa and lambda constant region genes.

The wave length is λ = 10-8  

The frequency is

f = s/ λ =

0.1 / 10-8  = 1/10 .10-8